OpenStudy (anonymous):

Find the slope-intercept form of the equation of the line passing through the given points. (-4, 0.6), (6, -2.4) y=

OpenStudy (anonymous):

i got answers for both, i'm not sure if they are both supposed to be the same or not after i use the point slope form.

OpenStudy (anonymous):

Find the gradient/slope of the line: (0.6 - -2.4)/ -4-6= -3/10 y= -3/10 x +c substituting point (-4, 0.6) 0.6 = -3/10 * -4 +c c=-0.6 equation y= -3/10 x-0.6